OpenStudy (xxkitookatxx):

Please help!!! What are the x-intercept(s) of the graph of y + 12 = x2 + x? A. (4, 0) and (3, 0) B. (4, 0) and (−3, 0) C. (−4, 0) and (−3, 0) D. (−4, 0) and (3, 0) I thought it was A. please tell me if im right.

OpenStudy (mathmale):

y + 12 = x2 + x would be better expressed as y + 12 = x^2 + x: the ^ symbol denotes exponentiation. If you want to find the x=intercepts, let y=0 and solve for x. Mind doing that and sharing your results? If you want the y-intercept, let x=0 and find y by solving the resulting equation.

OpenStudy (mathmale):

If you merely want to check your results, substitute the coordinates of the 2 points you've found back into the original equation. Is that equation now true? If so, you've got a solution. If not, it's back to the drawing board.

OpenStudy (qwertty123):

You can also use this online calculator. https://www.desmos.com/calculator @mathmale did a great job of explaining it!

OpenStudy (xxkitookatxx):

I know how to do it usually but im stumped on this!!!! Havent done it since 8th grade and i got all the other awnsers but no to this question. I did all the ways mt teacher taught me and i just for some reason i cant get to the awnser. IM guessing its not A?